### v1

# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

Abstract.
In this article, we consider eigenvector centrality for the nodes of a graph and study the robustness (and stability) of this popular centrality measure. For a given weighted graph $\mathcal{G}$ (both directed and undirected), we consider the associated weighted adiacency matrix $A$, which by definition is a non-negative matrix. Eigenvector centrality consists of ranking the elements of the graph according to the corresponding entries of the Perron eigenvector of $A$, which is associated with the positive eigenvalue with largest modulus.
An indicator of the robustness of eigenvector centrality consists in looking for a nearby perturbed graph $\widetilde{\mathcal{G}}$, with the same structure as $\mathcal{G}$ (i.e., with the same vertices and edges), but with a weighted adiacency matrix $\widetilde{A}$ such that the highest $m$ entries ($m\geq 2$) of the Perron eigenvector of $\widetilde{A}$ coalesce, making the ranking at the highest level ambiguous. To compute a solution to this matrix nearness problem, a nested iterative algorithm is proposed that makes use of a constrained gradient system of matrix differential equations (possibly on a low-rank manifold) in the inner iteration and a one-dimensional optimization of the perturbation size in the outer iteration.
The proposed algorithm produces the optimal perturbation (i.e., the one with smallest Frobenius norm) of the graph, which causes the looked-for coalescence, which is a measure of the sensitivity of the graph. The methodology is formulated in terms of graphs but applies to any nonnegative matrix, with potential applications in fields like population models, consensus dynamics, economics, etc.
Key words and phrases:
Discrete time linear switching system, spectral gap, Lyapunov exponent, joint spectral radius, dominant product, trajectories, asymptotic growth, finiteness property, algorithm, branch-and-bound.
Key words and phrases:
Eigenvector centrality; matrix nearness problem; structured eigenvalue optimization.
2020 Mathematics Subject Classification:
Primary 15A60, 05C50, 15A42; Secondary 47A55
Michele Benzi’s work was supported in part by MUR (Ministero dell’ Università e della Ricerca) through the PNRR MUR Project PE0000013-FAIR and through the PRIN Project 20227PCCKZ (“Low Rank Structures and Numerical Methods in Matrix and Tensor Computations and their Applications”).
Nicola Guglielmi acknowledges support from the Italian MUR within the PRIN 2022 Project “Advanced Numerical Methods for Time Dependent Parametric Partial Differential Equations with Applications”. He also thanks the support from the MIUR program for Departments of Excellence (project title of GSSI: Pattern analysis and engineering). Both authors acknowledge the MUR Pro3 joint project entitled “STANDS: Stability of Neural Dynamical Systems”. The authors are also affiliated to the INdAM-GNCS (Gruppo Nazionale di Calcolo Scientifico).
1. Introduction
In many fields, such as biology, social sciences, statistics, medicine and several others, ranking entities in a large data set is an important task. This leads to the concept of centrality measure for a network or a graph, which is the subject of this paper. High-rank (centrality) of a node in a graph indicates the importance of the node. Several centrality measures have been proposed in the literature; for a broad class of walk-based centrality measures see, e.g, [6, 4, 32]); here we consider eigencentrality, i.e., the one based on the Perron eigenvector, which was first introduced by Landau [29] and rediscovered by several authors since. For a strongly connected graph $\mathcal{G}$ with associated (possibly weighted) adjacency matrix $A$, this consists in ranking the nodes according to the entries of the Perron eigenvector of $A$, which is the positive eigenvector associated to the eigenvalue $\lambda>0$ of maximum modulus. As is well known, one the most important ranking schemes, the PageRank algorithm [7, 30], is based on a variant of this idea.
Given a directed or undirected weighted graph, we are interested in understanding how small perturbations may affect the ranking and which edges are more sensitive in order to produce certain changes in the ranking. We can distinguish two main motivations.
  • (i)
    Perturbing the weights as a robustness tool, or sensitivity analysis. Motivation: studying the reliability of the ranking in presence of uncertainties or variations in time of the weights, which in real networks are never known exactly.
  • (ii)
    Assessing the vulnerability of a network in the presence of deliberate attempts, for example by malicious, disgruntled individuals or by other interested agents, to manipulate the rankings by slightly perturbing the weights. E.g, in PageRank, one might want to manipulate the network in order to affect the rank of a product or of a political candidate.
A robustness indicator is useful in assessing the reliability of the measure. A very cheap and intuitively reasonable indicator of robustness consists of measuring the distance between the first and the $m$-th highest entries of the (suitably normalized) ranking eigenvector. However such a measure may be misleading, which means that even when this distance is small, this does not imply that the ranking is not robust, because - due to the structure of the graph - even a large modification of the weights of the graph might not be sufficient to annihilate the considered distance. For this reason we pose the problem in different terms and look for a closest graph $\widetilde{\mathcal{G}}$ to $\mathcal{G}$, which has in common with $\mathcal{G}$ the same sets of vertices and edges, but different weights, such that the largest entries of the Perron eigenvector coalesce. This problem translates into the search of a perturbed weighted adiacency matrix $\widetilde{A}$ to $A$ which preserves the sparsity pattern of $A$ and certain of its properties, and is characterized by colaescence of the first $m$-entries ($m\geq 2$) of the leading eigenvector.
In this way we define a structured robustness measure, whose computation will be discussed in this paper, together with the computation of the closest altered graph. It may be computationally expensive to be used as a general-purpose robustness indicator in comparison with simpler heuristic criteria, which are available without further computations, such as the above mentioned difference between the largest and $m$-th largest entry in the Perron eigenvector of $A$. But it has the advantage of a higher reliability. Moreover, cheap measures are not able to provide the nearby graph with ambiguous or changed ranking, which is a very important byproduct of the methodology we propose. The quantity considered in this paper is related to other structured robustness measures that arise in matrix and control theory, such as stability radii, passivity distances, distance to singularity, etc. [23, 21, 22, 2, 16, 18, 28].
A main objective of this paper is to show how the robustness measure can be computed through the explicit computation of a nearby graph with ambiguous ranking. The proposed algorithm is an iterative algorithm, where in each step the leading eigenvector of the weighted adjacency matrix of a graph with perturbed weights is computed. For a large sparse graph (where the number of edges leaving any vertex is bounded independently of $n$), these computations can typically be done with a more favorable complexity in the number $n$ of vertices with respect to a general purpose eigensolver. A feature of this algorithm in common with recent algorithms for eigenvalue optimization (such as the ones described in [16, 18, 1]) is a two-level methodology for matrix nearness problems, where in an inner iteration a gradient flow drives perturbations of a fixed norm to the original matrix into a (local) minimum of a nonnegative functional that depends on eigenvalues and/or eigenvectors, while in an outer iteration the perturbation size is optimized such that the functional becomes zero.
The paper is organized as follows. In Section 2 we briefly mention recent related work by other authors working on similar centrality manipulation problems. In Section 3 we recall basic concepts of eigenvector centrality measures. In Subsection 3.1 we introduce the radius of robustness ($\varrho_{m}$) for eigenvector centrality as the distance to the closest graph with coalescent highest $m$ entries in the associated Perron eigenvector. In Subsection 3.2 we formulate a methodology to compute the $\varrho_{m}$ and the nearby ambiguous graph. This requires the solution of a structured matrix nearness problem. For this we propose a two-level iterative method, which is based on a gradient system of matrix differential equations in an inner iteration and a one-dimensional optimization of the perturbation size in an outer iteration, and we discuss algorithmic aspects. In Section 4 we provide the necessary theoretical background on first order perturbation theory of eigenvalues and eigenvectors. In Section 5 we describe the complete computational procedure to compute the radius of robustness. In Section 6 we present the results of numerical experiments where the robustness measure discussed here is computed for some examples of graphs. In Section 7 we briefly address the stability of the Perron eigenvector. Finally, in Section 8 we draw our conclusions.
2. Related work
Related problems have been considered in the literature by several authors. Most relevant to our work are the papers [10, 13] and [33]; see also the references therein.
In [13] the authors consider the problem of perturbing a given stochastic matrix so that it has a prescribed stationary probability distribution, with the perturbation required to be minimal with respect to the matrix 1-norm and to preserve the sparsity pattern of the original matrix (besides of course, being such that the perturbed matrix remains stochastic). The solution is obtained by formulating the problem as a constrained linear optimization problem which is then solved by a suitable algorithm.
In [10], the authors consider two popular measures of centrality, Katz and PageRank (see, respectively, [32] and [30]). Their goal is to find the “minimal” perturbation (according to a certain matrix functional) of the weights that enforce a prescribed vector of centrality scores, while at the same time preserving the sparsity pattern. In particular, the problem of coalescing the top $m$ entries could be treated by the approach in [10], by enforcing a ranking vector in which the top $m$ entries take the same, prescribed value. On the other hand, PageRank and especially Katz centrality are not exactly the same as eigenvector centrality and, perhaps more important, the optimality measure used in [10] is not the Frobenius norm but a weighted sum of the (squared) Frobenius norm and the matrix 1-norm. Also, the approach in [10], based on quadratic programming, is quite different from the one we propose. Nevertheless, also in view of the close relationship existing between Katz and eigenvector centrality (see [6]), it is possible that the approach in [10] may be applicable, with some modifications, to the solution of the problem considered in the present paper.
Finally, we mention the interesting paper [33], which studies the broader problem of identifying sets of nodes in a network that can be used to “control” the eigenvector centrality scores of the remaining nodes by manipulating the weights of the corresponding outgoing links. In this paper it is shown how to identify minimal sets of nodes, called “controller nodes,” which can be used to obtain the desired eigenvector centrality ranking. This is clearly a different problem than the one considered in this paper, where we seek the perturbation of minimum Frobenius norm, but the approach in [33] may be combined with the one in our paper, since our method allows for the possibility of constraining the perturbations to a subset of the edges in the network (see the example in Section 5).
3. Robustness and vulnerability of eigenvector centrality
Let
$\mathcal{G}=(\mathcal{V},\mathcal{E},A)$
be a weighted graph (either directed or undirected) with vertex set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, edge set $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ and weighted adjacency matrix $A$. To each edge $(i,j)\in\mathcal{E}$ (from node $v_{j}$ to node $v_{i}$) is associated a positive weight $a_{ij}>0$; of course, for an undirected graph we have $a_{ji}=a_{ij}$ for all $i,j$. We set $a_{ij}=0$ for $(i,j)\notin\mathcal{E}$. Note that we exclude self-loops, hence $a_{ii}=0$ for all $i$. The weighted adjacency matrix of the graph is the non-negative matrix
$A=(a_{ij})\in\mathbb{R}^{n\times n}.$
When $G$ is undirected, $A$ is symmetric. To avoid a dependence of the measure on the scaling of the weights, we assume that $A$ is normalized with respect to the Frobenius norm,
(1) $\|A\|=\|A\|_{F}=1$
where in this article - unless specified differently - $\|\cdot\|=\|\cdot\|_{F}$.
Given a graph $\widetilde{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widetilde{A})$, with the same vertices and edges of $\mathcal{G}$, we define the distance between $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ by
(2) ${\rm dist}\bigl{(}\mathcal{G},\widetilde{\mathcal{G}}\bigr{)}=\sqrt{\sum\limits_{(i,j)\in\mathcal{E}}\left(a_{ij}-\widetilde{a}_{ij}\right)^{2}}=\|A-\widetilde{A}\|\,.$
We order the eigenvalues $\{\lambda_{i}\}$ of $A$ by non-increasing modulus so that $\lambda_{1}$ is the Perron eigenvalue of $A$, which is non-negative and equal to $\rho(A)$ (the spectral radius of $A$).
Assumption 3.1.
In this article we assume that the graph is strongly connected, which is equivalent to the irreducibility of $A$.
The following theorem is classic (see e.g. [25]).
Theorem 3.1 (Perron–Frobenius).
If $A\in\mathbb{R}^{n\times n}$ is nonnegative and irreducible then
  • (i)
    $\rho(A)>0$ is a simple eigenvalue of $A$;
  • (ii)
    there is a positive normalized eigenvector $\mathrm{v}$ such that $A\mathrm{v}=\rho(A)\mathrm{v}$ ($\mathrm{v}$ is called the Perron eigenvector of $A$).
Eigenvector centrality (see, e.g., [32]) is based on the ranking of the entries of the suitably normalized Perron eigenvector $\mathrm{v}$.
We indicate by $i_{1},i_{2},\ldots$ the indices of the entries of the Perron eigenvector $\mathrm{v}$ of $A$, ordered by non-decreasing magnitude. In the event that all the entries of $\mathrm{v}$ are distinct, the index $i_{1}$ corresponding to the largest entry in $\mathrm{v}$ identifies the most central node $i_{1}$ in $\mathcal{G}$, as measured by eigenvector centrality; $i_{2}$ identifies the second most central node, and so forth. Generally speaking, however, there could be many nodes with the same centrality score. An extreme case is when all the row sums of $A$ are equal, in which case eigenvector centrality is unable to provide a ranking of the nodes, every node being given the same centrality score.
As an illustrative example, for the following graph $\mathcal{G}$ on the left picture,
[…]
we have
$A=\left(\begin{array}[]{ccccc}0&14.9&6.7&3.9\\ 0&0&4.0&0\\ 0&5.2&0&14.1\\ 4.8&0&9.6&0\end{array}\right),\quad\mathrm{v}=\left(\begin{array}[]{c}0.5665\\ 0.1570\\ \mathbf{0.5844}\\ 0.5594\end{array}\right).\hskip 2.84544pt$
Hence, when ranked by (right) eigenvector centrality, $v_{3}$ is the most central node, $v_{1}$ the second most central one, and so forth.
In this case, the ranking provided by eigenvector centrality appears unambiguous. However, a relative perturbation of the weights of the graph of a few percents causes the two entries $\mathrm{v}_{1}$ and $\mathrm{v}_{3}$ to flip in magnitude, indicating a lack of robustness of the ranking.
We show this exhibiting by the nearby graph $\widetilde{\mathcal{G}}$ (on the right picture) with weighted adiacency matrix $\widetilde{A}$ and corresponding Perron eigenvector $\widetilde{\mathrm{v}}$:
$\widetilde{A}=\left(\begin{array}[]{rrrr}0&14.9&6.8&4.0\\ 0&0&4.1&0\\ 0&5.1&0&13.9\\ 4.7&0&9.5&0\end{array}\right),\qquad\widetilde{\mathrm{v}}=\left(\begin{array}[]{c}\mathbf{0.5774}\\ 0.1602\\ 0.5772\\ 0.5548\end{array}\right).$
Considering that in many applications the weights are known to an accuracy of only one or two digits, this example shows that indeed the ranking determined by eigenvector centrality may not always be reliable.
Generally speaking, the ranking of the vertices obtained by this approach becomes unreliable when one of the following occurrences is determined by a small perturbation of the weights:
  • (I)
    The coalescence of the entries $\mathrm{v}_{i_{1}},\mathrm{v}_{i_{2}},\ldots,\mathrm{v}_{i_{m}}$ for some index $m>1$.
  • (II)
    The coalescence of the eigenvalues $\lambda_{1}$ and $\lambda_{2}$.
Note that (II) can be ruled out if we make Assumption 3.1, i.e. in the irreducible case.
Indeed, if a small perturbation is able to coalesce the highest $m$ entries of the eigenvector $\mathrm{v}$, it means that the ranking for the highest positions is not robust, especially in the presence of uncertainties, rounding errors, or any other form of noise in the weights; in general this is a symptom of the fact that the highest $m$ vertices are difficult to rank using eigenvector centrality and it may therefore be better A similar issue arises when a small perturbation makes the adjacency matrix $A$ reducible (i.e., the graph is no longer strongly connected), in which case the Perron eigenvector may no longer be defined, rendering eigenvector centrality no longer usable. Note that issue (II) above falls under this scenario.
In this paper we direct our attention mainly to issue (I), and we are interested in computing the smallest perturbation of $A$ such that the colaescence of the highest entries in the Perron vector occurs. To this end, in the next subsection we propose a sharp measure which is aimed at characterizing the reliability of eigenvector centrality. As a by-product, the knowledge of this perturbation allows the possibility of modifying the ranking by small changes to the weights and may be used for this purpose. In Section 7 we briefly discuss issue (II).
3.1. Robustness of the ranking and nearest ambiguous graph
We consider the general case of a directed graph and look for the closest strongly connected graph $\widetilde{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widetilde{A})$ to $\mathcal{G}=(\mathcal{V},\mathcal{E},A)$, such that the $m$ highest entries of the Perron eigenvector coalesce. In order to preserve the strong connectivity we have to impose that the perturbed graph preserves the property $\widetilde{a}_{ij}>0$ for $(i,j)\in\mathcal{E}$. For computational purposes we replace it with the property
(3) $\widetilde{a}_{ij}\geq\delta\qquad\mbox{for}\quad(i,j)\in\mathcal{E},$
with $01\gg\delta>$ 0 a small positive tolerance.
We consider the following measure of robustness, which we call robustness radius,
$\varrho_{m}(\mathcal{G})=\>\inf_{\widetilde{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widetilde{\mathcal{G}}\right)=\inf_{\widetilde{A}}\|A-\widetilde{A}\|,$
under the condition that the $m$ largest entries of the Perron eigenvector of $\widetilde{A}$ coalesce and under the constraints $\widetilde{a}_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$ and $\widetilde{a}_{ij}=0$ for $(i,j)\notin\mathcal{E}$.
Note that we may choose to only perturb a subset of the edges, $\mathcal{E}^{\prime}\subset\mathcal{E}$, which would just change the constraint to $\widetilde{a}_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}^{\prime}$. For example, we may choose to perturb only the edges which either are subject to uncertainties or to changes in time.
Indeed it is possible (although rare) that no graph $\widetilde{\mathcal{G}}$ exists which fulfils the constraints, in which case we say that the distance is $\infty$. Moreover, when there exist graphs fulfilling the constraints, we have that the infimum is indeed a minimum, by compactness.
3.2. Outline of the computational approach
In this section we describe an approach to compute the robustness measure $\varrho_{m}(\mathcal{G})$ and the closest perturbed graph characterized by an ambiguous ranking of the most important $m$ vertices. It is convenient to write
$\widetilde{A}=A+\varepsilon E\quad\mbox{ with}\ \|E\|=1$
so that $\varepsilon$ is the norm of the perturbation to $A$.
Based on the fact that we are addressing the coalescence of $m$ distinct entries, we can make use of their average to obtain a concentration measure. For a given $\varepsilon>0$ and a given $E\in\mathbb{R}^{n\times n}$ with the same pattern of $A$ (and such that $A+\varepsilon E\geq 0$), we indicate by
$\mathrm{v}=\mathrm{v}(A+\varepsilon E)$
the Perron eigenvector of $\widetilde{A}$ and define the mean of its highest $m$ entries as
(4) $\Bigl{\langle}\mathrm{v}(A+\varepsilon E)\Bigr{\rangle}_{m}=\frac{1}{m}\sum_{j=1}^{m}\mathrm{v}_{i_{j}}\bigl{(}A+\varepsilon E\bigr{)},$
where $i_{1},i_{2},\ldots,i_{m}$ the indices of the $m$ largest entries of the Perron eigenvector $\mathrm{v}$ of $A$, ordered by non-decreasing magnitude.
Then we define the following functional of $E$ :
(5) $F_{\varepsilon}(E)$ $=$ $\frac{1}{2}\sum_{k=1}^{m}\Bigl{(}\mathrm{v}_{i_{k}}\bigl{(}A+\varepsilon E\bigr{)}-\Bigl{\langle}\mathrm{v}(A+\varepsilon E)\Bigr{\rangle}_{m}\Bigr{)}^{2}$
under the constraints of of preservation of the sparsity pattern of $A$, of unit Frobenius norm of $E$ and of non-negativity $A+\varepsilon E\geq 0$. This nonnegative functional measures the concentration of the highest $m$ entries of the Perron eigenvector $\mathrm{v}(A+\varepsilon E)$ with respect to their mean. Clearly the functional vanishes if and only if the $m$ highest entries of $\mathrm{v}(A+\varepsilon E)$ coalesce.
Our approach is summarized by the following two-level method:
  • Inner iteration: Given $\varepsilon>0$, we compute a real matrix $E=(e_{ij})\in\mathbb{R}^{n\times n}$ with the same sparsity pattern as $A$ (i.e., $e_{ij}=0$ if $a_{ij}=0$), with $E$ of unit Frobenius norm, with $(A+\varepsilon E)_{ij}\geq\delta$ (for $(i,j)\in\mathcal{E}$) such that the functional $F_{\varepsilon}(E)$ is minimized. The obtained minimizer is denoted by $E(\varepsilon)$.
  • Outer iteration: We compute the smallest value of $\varepsilon$ such that the $m$-th largest entries of the Perron eigenvector $\mathrm{v}\left(A+\varepsilon E(\varepsilon)\right)$ coalesce.
When there exists a solution to the problem $F_{\varepsilon}\left(E(\varepsilon)\right)=0$, in the outer iteration we compute the optimal $\varepsilon$, denoted $\widetilde{\varepsilon}_{m}$ and the corresponding extremal perturbation $E(\widetilde{\varepsilon}_{m})$,
(6) $\widetilde{\varepsilon}_{m}\longrightarrow\arg\min\limits_{\varepsilon>0}F_{\varepsilon}\left(E(\varepsilon)\right)=0$
by a combined Newton-bisection method.
Although in most interesting cases the equation $F_{\varepsilon}\left(E(\varepsilon)\right)=0$ has solutions, in general this is not assured. If a solution to $F_{\varepsilon}\left(E(\varepsilon)\right)=0$ exists then (6) yields
$\varrho_{m}(\mathcal{G})=\widetilde{\varepsilon}_{m}.$
Existence of local minima
An important issue concerns the existence of minima of (5). Let $\varepsilon>0$ be arbitrary but fixed. By the assumption of strong connectivity of the perturbed graph, we have that the eigenvalue $\lambda=\rho(A+\varepsilon E)$ is simple for all admissible perturbations $E$. This implies the continuity with respect to $E$ of the Perron eigenvector $v$ associated to $\lambda$, which implies the continuity of the functional $F_{\varepsilon}(E)$. Combining this and the fact that the set of considered perturbations is compact, a minimizer of (6) always exists. Since we make use of a gradient system approach to the problem, in general we cannot be sure to compute global optima, but only to globally converge to local optima, which means that the final computed matrix provides an upper bound to the solution of (6).
4. Derivatives of eigenvalues and eigenvectors
In this section we recall a few basic results concerning first order perturbation theory of real eigenvalues and eigenvectors of a real matrix, which is the relevant case for this paper.
4.1. First order perturbation theory for simple eigenvalues
The following standard perturbation result for eigenvalues is well-known; see e.g. [15, Theorem 1] and [27, Section II.1.1].
Theorem 4.1 (Derivative of simple eigenvalues).
Consider a continuously differentiable path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a continuous path of real simple eigenvalues of $M(t)$. Let $x(t)$ and $y(t)$ be left and right eigenvectors, respectively, of $M(t)$ to the eigenvalue $\lambda(t)$. Then, $x(t)^{\top}y(t)\neq 0$ for $t\in I$ and $\lambda$ is continuously differentiable on $I$ with the derivative (we denote ˙$\dot{\phantom{a}}=d/dt$)
(7) $\dot{\lambda}=\frac{x^{\top}\dot{M}y}{x^{\top}y}\,.$
Moreover, “continuously differentiable” can be replaced with “analytic” in the assumption and the conclusion.
Since we have $x(t)^{\top}y(t)\neq 0$, we can apply the normalization
(8) $\|x(t)\|=1,\ \ \|y(t)\|=1,\quad x(t)^{\top}y(t)\text{ is real and positive.}$.
4.2. First order perturbation theory for eigenvectors
For the derivative of eigenvectors we need the notion of group inverse (or reduced resolvent); see [31] as well as Kato ([27, Section I.5.3]).
Definition 4.1 (Group inverse).
Let $M\in{\mathbb{R}}^{n\times n}$ be a singular matrix with a simple zero eigenvalue. The group inverse (or reduced resolvent) of $M$ is the unique matrix $M^{\#}$ with
(9) $MM^{\#}=M^{\#}M,\qquad M^{\#}MM^{\#}=M^{\#},\quad\mbox{and}\quad MM^{\#}M=M.$
It is known from Meyer and Stewart ([31]) that if $M$ is normal, then its group inverse $M^{\#}$ is equal to the better known Moore–Penrose pseudoinverse $M^{\dagger}$. In general, the two pseudoinverses are not the same. They are, however, related by the following result, which is a special case of a more general result in [20, Appendix A] which is also verified directly.
Theorem 4.2 (Group inverse and Moore–Penrose pseudoinverse).
Suppose that the real matrix $M$ has the simple eigenvalue 00 with corresponding left and right real eigenvectors $x$ and $y$ of unit norm and such that $x^{\top}y>0$. Let $M^{\#}$ be the group inverse of $M$, and with $\kappa=1/(x^{\top}y)$ define the projection $\Pi={\rm I}-\kappa xy^{\top}$. The group inverse $M^{\#}$ of $M$ is related to the Moore–Penrose pseudoinverse $M^{\dagger}$ by
(10) $M^{\#}=\Pi M^{\dagger}\Pi.$
This result is used in the codes implemented by the authors for the computations involving the group inverse. Similar results are known in the Markov chain literature (see, e.g., [3]).
The group inverse appears in the following result on the derivative of eigenvectors, which is a variant of Theorem 2 of Meyer and Stewart ([31]).
Theorem 4.3 (Derivative of eigenvectors).
Consider a smooth path of square real matrices $M(t)$ for $t$ in an open interval $I$. Let $\lambda(t)$, $t\in I$, be a path of simple real eigenvalues of $M(t)$. Then, there exists a continuously differentiable path of associated left and right eigenvectors $x(t)$ and $y(t)$, $t\in I$, which are of unit norm with $x^{\top}(t)y(t)>0$ and satisfy the differential equations
(11) $\begin{array}[]{rcl}\dot{x}^{\top}\!\!\!&=&{}-x^{\top}\dot{M}M^{\#}+\left(x^{\top}\dot{M}M^{\#}x\right)x^{\top},\\[8.53581pt] \dot{y}&=&{}-M^{\#}\dot{M}y+\left(y^{\top}M^{\#}\dot{M}y\right)y,\end{array}$
where $M^{\#}(t)$ is the group inverse of $M(t)-\lambda(t)I$.
If $M$ is symmetric the previous simplify to $x=y$ and
(12) $\dot{x}^{\top}\!\!\!={}-x^{\top}\dot{M}M^{\dagger}$
where $M^{\dagger}(t)$ is the pseudoinverse of $M(t)-\lambda(t)I$.
Formulas (11) are interpreted as follows: the first term in the right-hand side represents the free derivative of the eigenvector, while the second term constrains the eigenvector to preserve the unit Euclidean norm with respect to $t$.
5. A methodology for computing the closest ambiguous graph and the robustness radius
In this section we describe the computational procedure to compute the robustness radius $\varrho_{m}(\mathcal{G})$. We indicate by
$\langle X,Y\rangle=\mathrm{trace}(X^{\top}Y)$
[…]
5.1. Inner iteration
We denote by $i_{1},i_{2},\ldots,i_{m}$ the indices of the $m$ largest entries of the Perron eigenvector $\mathrm{v}\left(A+\varepsilon E\right)$, ordered by non-decreasing magnitude. . Moreover, we denote the $i$-th unit versor by
[…]
For a set of edges $\mathcal{E}$, we define $P_{\mathcal{E}}$ as the orthogonal projection from $\mathbb{R}^{n\times n}$ onto the vector subspace of all real matrices having the sparsity pattern determined by $\mathcal{E}$: for $A=(a_{ij})$,
$P_{\mathcal{E}}(A)\big{|}_{ij}:=\begin{cases}a_{ij}\,,&\text{if }(i,j)\in\mathcal{E}\,,\\ 0\,,&\text{otherwise.}\end{cases}$
For a fixed given weight matrix $A$ and for $\varepsilon>0$, we call a matrix $E=(e_{ij})\in\mathbb{R}^{n\times n}$ $\varepsilon$-feasible if
  • (ia)
    $E$ is of unit Frobenius norm.
  • (ib)
    $E=P_{\mathcal{E}}(E)$.
  • (ic)
    $E$ is symmetric (only for undirected graphs).
  • (ii)
    $(A+\varepsilon E)_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$.
Further constraints
An extension to the case where only some of the weights are subject to perturbations is straightforward. For that, it is sufficient to define the subset $\mathcal{E}^{\prime}\subset\mathcal{E}$ of those arcs whose weights can be perturbed, and to replace everywhere the projection $P_{\mathcal{E}}$ by the analogous projection $P_{\mathcal{E}^{\prime}}$.
Similarly, if a few weights are related by a linear relationship, this can be easily included in the considered projection with respect to the Frobenius inner product.
5.2. Constrained gradient of the functional
Here we compute the gradient of the functional when restricted to matrices having the same sparsity pattern of $A$. The key idea is based on the fact that we are able to express the derivative of $F_{\varepsilon}(E)$ as $g(\dot{E})$ for a special function $g$. Using the simple fact that for vectors $b,c\in\mathbb{R}^{n}$ and a matrix $Z\in\mathbb{R}^{n\times n}$ one has
$b^{\top}Zc=\langle bc^{\top},Z\rangle,$
we are able to write (where $\dot{E}=\frac{d}{dt}E(t)$)
$g(\dot{E})=\langle Z_{\varepsilon}(E),\dot{E}\rangle$
with an appropriate matrix-valued function $Z_{\varepsilon}$, which - in a natural way - is interpreted as the constrained gradient of $F_{\varepsilon}$ (with respect to the Frobenius inner product). Projecting it onto the space of real matrices with sparsity pattern ${\mathcal{E}}$ (first constrain) and unit norm (second constrain), we obtain the sought fully constrained gradient.
Recalling (5), we have that $F_{\varepsilon}(E)=\frac{1}{2}\sum\limits_{k=1}^{m}\bigl{(}\mathrm{v}_{i_{k}}\left(A+\varepsilon E\right)-\left\langle\mathrm{v}(A+\varepsilon E)\right\rangle_{m}\bigr{)}^{2}$. After defining the auxiliary vector
$\bar{\rm u}=\frac{1}{m}\sum_{j=1}^{m}{\rm u}_{i_{j}}$
we are able to state the following result.
Lemma 5.1 (Fully constrained gradient of the functional).
Consider a smooth path $E(t)$ of $\varepsilon$-feasible matrices, denote by $\lambda(t)=\rho\left(A+\varepsilon E(t)\right)$ the Perron eigenvalue and $\mathrm{v}(t)=\mathrm{v}\left(A+\varepsilon E(t)\right)$ the associated Perron eigenvector. Let $i_{1},i_{2},\ldots$ the indices of the entries of $\mathrm{v}(t)$, ordered by non-decreasing magnitude. We have (omitting the omnipresent dependence on $t$)
(13) $\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=\eta\Big{\langle}\widehat{G}_{\varepsilon}(E),\dot{E}\Big{\rangle},\qquad\widehat{G}_{\varepsilon}(E)=G_{\varepsilon}(E)-\big{\langle}G_{\varepsilon}(E),E\big{\rangle}E,$
where $\eta>0$ is a suitable scaling factor and $\widehat{G}_{\varepsilon}(E)$ is the fully constrained gradient of $F_{\varepsilon}$ with
(14) $G_{\varepsilon}(E)=P_{\mathcal{E}}\left(\sum_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\Bigl{(}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)(A_{\varepsilon}^{\#})^{\top}\mathrm{v}\mathrm{v}^{\top}-(A_{\varepsilon}^{\#})^{\top}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}\right)$
the structure constrained gradient, having the sparsity pattern determined by the set of edges $\mathcal{E}$.
Proof.
By (5) we get
$\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=\sum_{k=1}^{m}\bigl{(}\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\bigr{)}\bigl{(}\dot{\mathrm{v}}_{i_{k}}-\langle\dot{\mathrm{v}}\rangle_{m}\bigr{)}.$
Lemma 4.1 applied to $A+\varepsilon E(t)$ yields
(15) $\dot{\mathrm{v}}_{i_{k}}-\langle\dot{\mathrm{v}}\rangle_{m}$ $=$ $\Bigl{(}{\rm u}_{i_{k}}-\frac{1}{m}\sum_{j=1}^{m}{\rm u}_{i_{j}}\Bigr{)}^{\top}\dot{\mathrm{v}}=\varepsilon\left({\rm u}_{i_{k}}-\bar{\rm u}\right)^{\top}\left(-A_{\varepsilon}^{\#}\dot{E}\mathrm{v}+\left(\mathrm{v}^{\top}A_{\varepsilon}^{\#}\dot{E}\mathrm{v}\right)\mathrm{v}\right)$
$=$ $\varepsilon\Bigl{(}\big{\langle}-(A_{\varepsilon}^{\#})^{\top}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top},\dot{E}\big{\rangle}+\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\big{\langle}(A_{\varepsilon}^{\#})^{\top}\mathrm{v}\mathrm{v}^{\top},\dot{E}\big{\rangle}\Bigr{)}$
where $A_{\varepsilon}^{\#}(t)$ is the group inverse of $A+\varepsilon E(t)-\lambda(t)I$ and $\mathrm{v}(t)$ is the Perron eigenvector of $A+\varepsilon E(t)$, normalized with $\|\mathrm{v}(t)\|=1$. Hence we can write
$\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=\varepsilon\big{\langle}R,\dot{E}\big{\rangle}$
with the free gradient
$R=\sum_{k=1}^{m}\bigl{(}\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\bigr{)}\Biggl{(}\bigl{(}\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\bigr{)}(A_{\varepsilon}^{\#})^{\top}\mathrm{v}\mathrm{v}^{\top}-(A_{\varepsilon}^{\#})^{\top}\bigl{(}{\rm u}_{i_{k}}-\bar{\rm u}\bigr{)}\mathrm{v}^{\top}\Biggr{)}.$
Using the property that $P_{\mathcal{E}}(\dot{E})=\dot{E}$, we simply write
$\big{\langle}R,\dot{E}\big{\rangle}=\big{\langle}P_{\mathcal{E}}(R)+\left(R-P_{\mathcal{E}}(R)\right),\dot{E}\big{\rangle}$
and observe that
$\big{\langle}R-P_{\mathcal{E}}(R),\dot{E}\big{\rangle}=\big{\langle}R-P_{\mathcal{E}}(R),P_{\mathcal{E}}(\dot{E})\big{\rangle}=0,$
This gives the structure-constrained gradient $G_{\varepsilon}(E)=P_{\mathcal{E}}(R)$ (see (14)).
In order to include the further norm constraint (ia) we project the gradient onto the tangent space to the set $\mathcal{S}_{1}=\{M\in\mathbb{R}^{n\times n}\ |\ \|M\|=1\}$ at $E$. To comply with the constraint ${\|E(t)\|^{2}=1}$, we must have
(16) $0=\frac{1}{2}\,\frac{d}{dt}\|E(t)\|^{2}=\langle E(t),\dot{E}(t)\rangle.$
We are thus led to the constrained optimization problem $\arg\min\limits_{\langle E,Z\rangle=0}\ \langle G,Z\rangle$ (where $G=G_{\varepsilon}(E)$). In order to obtain a unique solution we also impose that the solution (which represents a direction) has unit norm. Thus we have to solve the problem
(17) $Z_{\star}$ $=$ $\arg\min_{\|Z\|=1,\,\langle E,Z\rangle=0}\ \langle G,Z\rangle.$
Its solution is given by
(18) $\mu Z_{\star}$ $=G-\langle E,G\rangle\,E,$
where $\mu$ is the Frobenius norm of the matrix on the right-hand side. The solution is unique if $G$ is not a multiple of $E$. The result follows on noting that the real inner product with a given vector (which here is a matrix) is maximized over a subspace by orthogonally projecting the vector onto that subspace. The expression in (18) is the orthogonal projection of $G$ onto the orthogonal complement of the span of $E$, which is the tangent space at $E$ of the manifold of matrices of unit Frobenius norm. Setting $\widehat{G}_{\varepsilon}(E)$ as the right-hand-side of (18) gives (13). This concludes the proof. ∎
Constrained gradient system
In order to minimize the functional $F_{\varepsilon}$ over the space of $\varepsilon$-feasible matrices it is natural to consider the associated gradient system
(19) $\dot{E}=-\widehat{G}_{\varepsilon}(E)=-G_{\varepsilon}(E)+\gamma E\quad\hbox{ with }\quad\gamma=\langle E,G_{\varepsilon}(E)\rangle.$
It may happen that along the solution trajectory of (19) some positive entry of $A+\varepsilon E$ may reach the small positive value $\delta$. In our experiments we never observed such an occurrence, although this is a potential situation. In subsection 5.4 we comment on how to deal with this further constraint.
Thus we treat the constraint (ii) (which is a relaxation of non-negativity) by penalization. Omitting for now penalization (which we shall discuss later), we analyze in the next section the peculiar properties of the system of ODEs (19).
5.3. Monotonicity and stationary points
The following monotonicity result follows directly from the construction of the gradient system (19) (for the moment we assume that the non-negativity constraint does not activate so that penalization is not necessary, which is what we observe in all the tests we have done).
Theorem 5.1.
Let $E(t)$ of unit Frobenius norm satisfy the differential equation (19) with $G_{\varepsilon}(E)$ as in (14). Then, the functional $F_{\varepsilon}(E(t))$ decreases monotonically with respect to $t$:
$\dot{F}_{\varepsilon}(E(t))\leq 0.$
Furthermore, the following result characterizes stationary points.
Theorem 5.2.
The following statements are equivalent along solutions of (19):
  1. (1)
    ${\dot{F}}_{\varepsilon}=0$.
  2. (2)
    $\dot{E}=0$.
  3. (3)
    $E$ is a real multiple of $G_{\varepsilon}(E)$.
Proof.
Using Lemma 5.1 and (19) we obtain, with $G=G_{\varepsilon}(E)$,
(20) $\frac{1}{\varepsilon}\,\dot{F}_{\varepsilon}=\langle G,\dot{E}\rangle=\langle G,-G+\gamma E\rangle=-\|G\|^{2}+\langle G,E\rangle^{2}\leq 0.$
By (20), we have that
$\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=0\qquad\Longrightarrow\qquad\mbox{$E$ is a multiple of $G$}.$
Then $\langle G,\dot{E}\rangle$ with $\dot{E}=-G+\gamma E$ implies $\dot{E}=0$.
Clearly, 3. implies 2., which implies 1. Since the Cauchy–Schwarz inequality in (20) is strict unless $E$ is a multiple of $G$, we obtain that 1. implies 3. ∎
5.4. Underlying low-rank property of extremizers
For a sparse graph we have a number of ODEs which is given by the nonzero entries of the sparsity pattern associated to the graph. However, for a dense graph we may have up to $n^{2}$ ODEs. Following the lines of [19], we are able to exploit the underlying low-rank structure of the solutions of the optimization problem we consider. By Theorem 5.1 we have that the stationary points which we aim to compute are proportional to the gradient $G$, which - as we have seen - is the projection of a rank-$1$ matrix. Proceeding similarly to [19], by writing $E=P_{\mathcal{E}}(H)$ for a rank-$1$ matrix $H$, we are able to obtain a new system of ODEs for $H$, whose dynamics is constrained to the manifold of rank-$1$ matrices. We can prove that the stationary points of the two systems of ODEs are in a one-to-one correspondence and also that the system of ODEs for $H$, even if it loses the gradient system structure, converges - at least locally - to the stationary points. We omit a full derivation for sake of conciseness and refer the reader to [19] for details.
5.5. Non-negativity constraint
A further constraint on $E(t)$ is relaxed non-negativity, which means $(A+\varepsilon E(t))_{ij}\geq\delta$ (with $\delta$ a very small tolerance) for all $t$.
One possibility would be to consider the Karush-Kuhn-Tucker conditions by managing active constraints and integrating a piecewise smooth system of differential equations. However, due to the fact that in our experiments this situation has never occurred, we expect to observe it only exceptionally. For this reason we treat this situation by simply adding a suitable penalty term to the functional $F_{\varepsilon}(E)$,
$F_{\varepsilon}^{c}(E)=F_{\varepsilon}(E)+\frac{c}{\varepsilon}\,P_{\varepsilon}(E)$
for increasing $c\gg\varepsilon$, with
(21) $P_{\varepsilon}(E)=\frac{1}{2}\sum\limits_{(i,j)\in\mathcal{E}}\left(a_{ij}+\varepsilon e_{ij}-\delta\right)_{-}^{2}$
where $(p)_{-}=\min\left(p,0\right)$. Differentiating $P_{\varepsilon}$ yields
$\nabla P_{\varepsilon}(E)=\varepsilon\sum\limits_{(i,j)\in\mathcal{E},a_{ij}+\varepsilon e_{ij}<\delta}\left(a_{ij}+\varepsilon e_{ij}-\delta\right)_{-}$
which has to be included in the gradient system (19). The modified matrix-valued ODE is
(22) $\dot{E}=-G_{\varepsilon}^{c}(E)+\gamma E\quad\hbox{ with }\quad\gamma=\langle E,G_{\varepsilon}^{c}(E)\rangle$
where
(23) $G_{\varepsilon}^{c}(E)=G_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E).$
Whenever a trajectory $E(t)$ determines a negative entry of $A+\varepsilon E(t)$, the gradient system tends to recover non-negativity, with an effect which is proportional to $c$.
Similarly to Theorems 5.1 and 5.2 we have the following results, whose proofs are omitted since they are analogous to those of previous theorems.
Theorem 5.3 (Monotonicity of the penalized functional).
Let $E(t)$ of unit Frobenius norm satisfy the differential equation (22) with arbitrary but fixed $c>0$ and $G^{c}_{\varepsilon}(E)$ given by (23). Then, the functional $F_{\varepsilon}^{c}(E(t))$ decreases monotonically with respect to $t$: $\dot{F}^{c}_{\varepsilon}(E(t))\leq 0.$
Theorem 5.4 (Stationary points of the penalized functionl).
The following statements are equivalent along solutions of (22):
  1. (1)
    ${\dot{F}}_{\varepsilon}^{c}=0$.
  2. (2)
    $\dot{E}=0$.
  3. (3)
    $E$ is a real multiple of $G_{\varepsilon}^{c}(E)$.
5.6. Undirected graphs
Undirected graphs lead to some significant simplifications. In this case we have that $A_{\varepsilon}^{\#}=A_{\varepsilon}^{\dagger}$, the Moore-Penrose pseudoinverse, which is symmetric. We define for $A\in\mathbb{R}^{n\times n}$,
$\mathrm{Sym}(A)\big{|}_{ij}:=\frac{a_{ij}+a_{ji}}{2}$
By the well-known property $\mathrm{v}^{\top}A_{\varepsilon}^{\dagger}=0$, and including the symmetry constraint (ic), the gradient (14) simplifies to
(24) $G_{\varepsilon}(E)$ $=$ ${}-P_{\mathcal{E}}\left(\sum_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\mathrm{Sym}\Bigl{(}A_{\varepsilon}^{\dagger}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}\right)$
$=$ ${}-P_{\mathcal{E}}\left(\sum_{k=1}^{m}c_{k}\mathrm{Sym}\Bigl{(}a_{k}\mathrm{v}^{\top}\Bigr{)}\right),$
where we have defined
$a_{k}=A_{\varepsilon}^{\dagger}b_{k}=(A+\varepsilon E-\lambda I)^{\dagger}b_{k},\qquad\mbox{with}\qquad b_{k}=\left({\rm u}_{i_{k}}-\bar{\rm u}\right),\qquad c_{k}=\bigl{(}\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\bigr{)},$
which may be computed by solving the nonsingular linear system
(25) $\begin{pmatrix}A+\varepsilon E-\lambda I&\mathrm{v}\\ \mathrm{v}^{\top}&0\end{pmatrix}\begin{pmatrix}a_{k}\\ \mu\end{pmatrix}=\begin{pmatrix}b_{k}\\ 0\end{pmatrix}.$
Note that the submatrix $A+\varepsilon E-\lambda I$ above is negative semidefinite and singular. It is straightforward to check that the linear system (25) is nonsingular if and only if the orthogonal complement of $\mathrm{v}$ and the kernel of $A+\epsilon E-\lambda I$ have trivial intersection. Since $\text{Ker}(A+\varepsilon E-\lambda I)=\text{Span}\,\{\mathrm{v}\}$ and obviously $\mathrm{v}^{\top}\mathrm{v}\neq 0$, the condition is satisfied and the system (25) has a unique solution.
The computational cost of computing $G_{\varepsilon}(E)$ consists of computing the Perron eigenvector of the matrix $A+\varepsilon E(t)$ for $t$ a suitable point on the mesh of integration and in solving – at every step – the linear system (25). The Perron eigenvector can be computed efficiently, for example using Lanczos’ method; if the Perron root is well separated by the rest of the spectrum, even the power method will converge quickly. The solution of the linear system (25) merits a more detailed discussion.
Solving the linear system (25)
In the considered undirected case, system (25) is symmetric and indefinite, with one positive and $n$ negative eigenvalues. For small and moderate values of $n$, the solution can be obtained by means of some variant of the LDLT factorization, see [14]. For large and sparse problems, an efficient way to solve (25) is by means of the Minimal Residual (MinRes) iteration [34], possibly combined with a suitable symmetric positive definite preconditioner [5]. Alternatively, system (25) can be replaced by the equivalent one
(26) $\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&\mathrm{v}\\ \mathrm{v}^{\top}&0\end{pmatrix}\begin{pmatrix}a_{k}\\ \mu\end{pmatrix}=\begin{pmatrix}b_{k}\\ 0\end{pmatrix}.$
Note that the submatrix $A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}$ is negative definite (nonsingular). This matrix is completely full, since $\mathrm{v}>0$, and it should not be formed explicitly. System (26) can be solved by means of the block LU factorization
(27) $\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&\mathrm{v}\\ \mathrm{v}^{\top}&0\end{pmatrix}=\begin{pmatrix}A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}&0\\ \mathrm{v}^{\top}&-s\end{pmatrix}\begin{pmatrix}I&(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}\\ 0&1\end{pmatrix}\,,$
where we have set $s=\mathrm{v}^{\top}(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}$. It follows that the solution of (26), and therefore of (25), is given by
$a_{k}$ $=$ $(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}b_{k}-\mu(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}\,,$
$\mu$ $=$ $\frac{\mathrm{v}^{\top}(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}b_{k}}{\mathrm{v}^{\top}(A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top})^{-1}\mathrm{v}}\,,$
showing that this approach requires the solution of two linear systems with the same coefficient matrix $A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}$ and right-hand sides $b_{k}$ and $\mathrm{v}$, respectively. These two linear systems can be solved by means of the conjugate gradient (CG) method, possibly combined with a preconditioner. Clearly, the CG method is well suited since matrix-vector products involving $A+\varepsilon E-\lambda I-\mathrm{v}\mathrm{v}^{\top}$ can be computed without forming the coefficient matrix explicitly; for sparse $A$, the cost of each matrix-vector product is linear in $n$. The choice of preconditioner is clearly problem dependent; note, however, that as the outer iteration converges, the coefficient matrix of the linear systems varies slowly from one step to the next, making it possible to reuse the same preconditioner for several steps.
Projected rank-$2$ properties of extremizers
We let
$L_{\varepsilon}(E)=\left(\sum_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\mathrm{Sym}\Bigl{(}A_{\varepsilon}^{\dagger}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}\right)$
and $G_{\varepsilon}(E)={}-P_{\mathcal{E}}\left(L_{\varepsilon}(E)\right)$. Note that $A_{\varepsilon}^{\dagger}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}$ is a rank-$1$ nilpotent matrix. We let
$g_{i_{j}}=A_{\varepsilon}^{\dagger}{\rm u}_{i_{j}},\qquad\mbox{the $i_{j}$-th column of}\ A_{\varepsilon}^{\dagger}$
and
$\tau_{i_{k}}=\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m},\qquad\qquad a_{i_{k}}=g_{i_{k}}-\langle g\rangle_{m},\quad\mbox{with}\quad\langle g\rangle_{m}=\frac{1}{m}\sum_{j=1}^{m}g_{i_{j}}.$
This notation allows us to rewrite
$L_{\varepsilon}(E)=\mathrm{Sym}\left(a\mathrm{v}^{\top}\right),\qquad a=\sum_{k=1}^{m}\tau_{i_{k}}a_{i_{k}},$
which is a rank-$2$ symmetric matrix.
If $(i,j)\in\mathcal{E}$ for all $i=1,\ldots,n$ and $j=1,\ldots,n$ (which includes possible self-loops), we have that the pattern $\mathcal{E}$ is full, so that $P_{\mathcal{E}}$ is the identity projector. Thus $G_{\varepsilon}(E)=-L_{\varepsilon}(E)$, which means that the stationary points (if the penalty function is inactive) are rank-$2$.
In the general case, stationary points are projections onto $\mathcal{E}$ of rank-$2$ symmetric matrices.
5.7. An illustrative example of an undirected graph
[…]
Figure 1. Example of symmetric weighted graph
Consider the simple undirected graph with $9$ vertices of Figure 1. From a topological point of view one could argue that the vertex $v_{4}$ is characterized by the highest centrality score; however, since the graph is weighted this is not true in general.
In order to associate to the graph a scaled adiacency matrix $A$ with unit Frobenius norm, we multiply the adiacency matrix whose weights are indicated in Figure 1 by the normalizing factor
$\beta\approx 1/0.9974\approx 1.0026.$
We have $\lambda=\rho(A)\approx 0.5559\ldots$ and the associated Perron eigenvector (to $4$ digits)
$\mathrm{v}\approx\left(\begin{array}[]{ccccccccc}\mathbf{0.4844}&0.2712&0.2602&\mathbf{0.4553}&0.2154&0.2433&0.2941&0.2082&0.4259\end{array}\right)^{\top}.$
The ranking of nodes $\mathrm{v}_{1}$ and $\mathrm{v}_{4}$ can be interchanged by small changes in the weights. This is illustrated in Figure 2. Choosing a suitable small perturbation (of norm $0.0280$) of the normalized adiacency matrix $A$, we obtain the modified Perron eigenvector $\tilde{\mathrm{v}}$
$\tilde{\mathrm{v}}=\left(\begin{array}[]{ccccccccc}\mathbf{0.4754}&0.2699&0.2722&\mathbf{0.4758}&0.2277&0.2543&0.2839&0.2035&0.4027\end{array}\right)^{\top}.$
where now $\mathrm{v}_{4}$ precedes $\mathrm{v}_{1}$.
Note that, by a simple continuity argument of the Perron eigenvector, there exists a perturbation of the weights for which $\tilde{\mathrm{v}}_{1}$ and $\tilde{\mathrm{v}}_{4}$ have exactly the same value.
[…]
Figure 2. Perturbed graph of Figure 1 (weights of modified graph appear in bold on each edge).
We are interested in the robustness of the ranking for the first $m=2$ positions, corresponding to $i_{1}=1$ and $i_{2}=4$. The gap between the $1$-st and $4$-th entries is $\mathrm{v}_{1}-\mathrm{v}_{4}=0.0291\ldots$
According to (24), we compute the structure-constrained gradient $G_{0}$, normalized to have Frobenius norm equal to $1$; it is given by
$G_{0}=\left(\begin{array}[]{rrrrrrrrr}0&\!0.336&\!0&\!-0.113&\!0&\!0&\!0&\!0&\!0.645\\ 0.336&\!0&\!-0.029&\!0&\!0&\!0&\!0&\!0&\!0\\ 0&\!-0.029&\!0&\!-0.412&\!0&\!0&\!0&\!0&\!0\\ -0.113&\!0&\!-0.412&\!0&\!-0.419&\!-0.438&\!0&\!0&\!0\\ 0&\!0&\!0&\!-0.418&\!0&\!0&\!0&\!-0.079&\!0\\ 0&\!0&\!0&\!-0.438&\!0&\!0&\!-0.074&\!0&\!0\\ 0&\!0&\!0&\!0&\!0&\!-0.074&\!0&\!0&\!0.188\\ 0&\!0&\!0&\!0&\!-0.073&\!0&\!0&\!0&\!0.109\\ 0.645&\!0&\!0&\!0&\!0&\!0&\!0.188&\!0.109&\!0\\ \end{array}\normalsize\right).$
This indicates that a perturbation on the weights $a_{19},a_{12},a_{34},a_{45}$ and $a_{46}$ has the highest influence on the gap $\mathrm{v}_{1}-\mathrm{v}_{4}$.
Modifying the full pattern of $A$
The choice $\varepsilon=0.029$ with the natural initial value for the ODE (19) given by $E(0)=-G_{0}$ turns out to give interesting results. After integration of the ODE (19), we get the Perron eigenvector of the (optimally) perturbed matrix $A+\varepsilon E(\varepsilon)$,
$\widetilde{\mathrm{v}}=\small\left(\begin{array}[]{ccccccccc}\mathbf{0.4742}&0.2695&0.2722&\mathbf{0.4753}&0.2284&0.2551&0.2848&0.2044&0.4029\\ \end{array}\right)^{\top}\normalsize$
which shows an inversion of the ranking of the leading nodes, in fact
(28) $\widetilde{\mathrm{v}}_{1}-\widetilde{\mathrm{v}}_{4}=-0.0011\ldots\qquad\Longrightarrow\qquad\varrho_{2}(\mathcal{G})<0.029,$
where we recall that $\varrho_{2}(\mathcal{G})=\>\min_{\widetilde{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widetilde{\mathcal{G}}\right)=\min_{\widetilde{A}}\|A-\widetilde{A}\|,$ under the condition that the $2$ largest entries of the Perron eigenvector of $\widetilde{A}$ coalesce and under the constraints $\widetilde{a}_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$ and $\widetilde{a}_{ij}=0$ for $(i,j)\notin\mathcal{E}$. It is also interesting to observe that the gap between $\widetilde{\mathrm{v}}_{4}$ and $\widetilde{\mathrm{v}}_{9}$ on the other hand has increased.
A further tuning of the parameter $\varepsilon$ in the outer iteration provides the value $\varepsilon^{*}\approx 0.0279064$ as the one leading to coalescence (from a numerical point of view, coalescence was declared as soon as the condition $|\widetilde{\mathrm{v}}_{1}-\widetilde{\mathrm{v}}_{4}|\leq{\rm tol}$ with ${\rm tol}=10^{-5}$ was satisfied). We obtain in fact
$\widetilde{\mathrm{v}}\approx\left(\begin{array}[]{ccccccccc}\mathbf{0.4745}&0.2696&0.2717&\mathbf{0.4745}&0.2279&0.2546&0.2851&0.2045&0.4038\end{array}\right)^{\top}.$
Based on an heuristic argument, one may consider a perturbation $E$ whose only nonzero entries are $e_{12}=e_{21}=e_{14}=e_{41}=e_{19}=e_{91}=-1/\sqrt{6}$ (which guarantess the unit Frobenius norm of $E$), aiming to decrease the weights on the edges which connect the vertex $v_{1}$, which has highest centrality score, to the rest of the graph. With a perturbation again of magnitude $\varepsilon=0.029$, this would give the Perron eigenvector of $A+\varepsilon E$,
$\overline{\mathrm{v}}\approx\small\left(\begin{array}[]{ccccccccc}\mathbf{0.4689}&0.2640&0.2624&\mathbf{0.4543}&0.2229&0.2523&0.3047&0.2161&0.4270\end{array}\right)^{\top}\normalsize$
which preserves the original unperturbed ranking and is characterized by a gap of $0.0146$, significantly different (and with the oppsite sign) from the one in (28).
[…]
Figure 3. Modified graph of Figure 1 acting only on a subset of the edges (modified weights on each edge appear in bold)
Modifying only a subpattern of $A$
As we have mentioned, it is possible to consider, instead of the whole set of edges, a subset $\tilde{\mathcal{E}}\subset\mathcal{E}$ and proceed exactly as discussed by simply replacing the projection $P_{\mathcal{E}}$ with $P_{\tilde{\mathcal{E}}}$. As an example we consider again the simple undirected graph with $9$ vertices of Figure 1 and modify only the edges not incident to the two nodes $v_{1}$ and $v_{4}$. The results so obtained are illustrated in Figure 3. We compute the Perron eigenvector $\mathrm{v}$
$\mathrm{v}\approx\left(\begin{array}[]{ccccccccc}\mathbf{0.4913}&0.2916&0.2856&\mathbf{0.4913}&0.2406&0.2608&0.2433&0.1894&0.3601\end{array}\right)^{\top}.$
and the distance $\varepsilon^{*}=0.1407018$,
5.8. Numerical integration of the gradient system
We describe here the $k$-th step of numerical integration of (22). Starting from $E_{k}\approx E(t_{k})$ feasible matrix of unit norm, we apply a step of length $h$ of the explicit Euler method,
(29) $\tilde{E}_{k+1}=E_{k}+h\Bigl{(}{}-G_{\varepsilon}^{c}(E_{k})+\gamma_{k}E_{k}\Bigr{)}\quad\hbox{ with }\quad\gamma_{k}=\Big{\langle}E_{k},G_{\varepsilon}^{c}(E_{k})\Big{\rangle}$
where $G_{\varepsilon}^{c}$ has been computed according to (23).
We are not interested in following the exact solution of (22) but mainly to decrease the functional $F_{\varepsilon}^{c}$. This is equivalent to a gradient descent algorithm with line-search. According to this, for the stepsize selection we require that
(30) $F_{\varepsilon}^{c}(\tilde{E}_{k+1})<F_{\varepsilon}^{c}(E_{k}).$
Moreover, we apply an Armijo-like rule. For Equation (22) we have
$\frac{d}{dt}F^{c}_{\varepsilon}(E(t))\Big{|}_{t=t_{k}}=-\varepsilon\Bigl{(}\|G_{\varepsilon}^{c}(E_{k})\|^{2}-\langle G_{\varepsilon}^{c}(E_{k}),E_{k}\rangle^{2}\Bigr{)}:=-g_{k}\leq 0.$
We accept the result of the step with step size $h$ and set $E_{k+1}=\tilde{E}_{k+1}$ if (30) holds, otherwise we reduce the stepsize to $h/\theta$ for some fixed $\theta>1$, and repeat it. Moreover, if
(31) $F_{\varepsilon}^{c}(\tilde{E}_{k+1})\geq F_{\varepsilon}^{c}(E_{k})-(h/\theta)g_{k},$
then we reduce the step size for the next step to $h/\theta$. If the step size has not been reduced in the previous step, we then increase the step-length to be used in $[t_{k+1},t_{k+2}]$.
The computational cost of a single step requires the computation of the Perron eigenvector and the solution of the linear systems involving either the group inverse or, in the undirected case, the Moore-Penrose pseudoinverse (see (25)). In the undirected case we have to compute precisely $m$ columns of the Moore-Penrose pseudoinverse.
5.9. Outer iteration
Let $E(\varepsilon)$ denote the minimizer of the functional $F_{\varepsilon}^{c}$. Generically we expect that for a given perturbation size $\varepsilon<\widetilde{\varepsilon}_{m}$, $E(\varepsilon)$ is smooth, which means that the spectral radius of $A+\varepsilon E$ is a simple eigenvalue. If so, then $f_{c}(\varepsilon):=F_{\varepsilon}^{c}(E(\varepsilon))$ is a smooth function of $\varepsilon$ and we can exploit its regularity to obtain a fast iterative method to converge to $\widetilde{\varepsilon}_{m}$ from the left. Otherwise we can use a bisection technique to approach $\widetilde{\varepsilon}_{m}$.
An example is provided in Figure 4. Since the penalty term is never active the value of $c$ is not relevant here.
The following result provides an inexpensive formula for the computation of the derivative of $f_{c}(\varepsilon)=F_{\varepsilon}^{c}(E(\varepsilon))$, which will be useful in the construction of the outer iteration of the method.
Assumption 5.1.
We assume that $E(\varepsilon)$ is a smooth function of $\varepsilon$ in some interval.
We then have the following result.
Lemma 5.2.
Under Assumption 5.1, the function $f_{c}(\varepsilon)=F_{\varepsilon}^{c}(E(\varepsilon))$ is differentiable and its derivative equals (with ${\phantom{a}{}^{\prime}}=d/d\varepsilon$)
(32) $f_{c}^{\prime}(\varepsilon)=-\|G_{\varepsilon}^{c}(E(\varepsilon))\|$
Proof.
The conservation of $\|E(\varepsilon)\|=1$ implies
$\Big{\langle}E(\varepsilon),E^{\prime}(\varepsilon)\Big{\rangle}=0.$
Differentiating $f_{c}(\varepsilon)=F_{\varepsilon}^{c}\bigl{(}E(\varepsilon)\bigr{)}$ with respect to $\varepsilon$ we obtain
(33) $\hskip-25.60747ptf_{c}^{\prime}(\varepsilon)=\Bigl{\langle}G_{\varepsilon}^{c}(E(\varepsilon)),E(\varepsilon)+\varepsilon E^{\prime}(\varepsilon)\Bigr{\rangle}=\Bigl{\langle}G_{\varepsilon}^{c}(E(\varepsilon)),E(\varepsilon)\Bigr{\rangle}.$
By means of the property no. 3. of minimizers, stated in Theorem 5.4, we can conclude that $E(\varepsilon)\propto G_{\varepsilon}^{c}(E(\varepsilon))$, which yields formula (32). ∎
For $\varepsilon<\widetilde{\varepsilon}_{m}$, we make use of the standard Newton iteration
(34) $\varepsilon_{1}=\varepsilon_{o}-\frac{f_{c}(\varepsilon_{o})}{f_{c}^{\prime}(\varepsilon_{o})},$
to get an updated value $\varepsilon_{1}$ from $\varepsilon_{o}$.
In a numerical code it is convenient to couple the Newton iteration (34) with the bisection method, to increase its robustness. We make use of a tolerance which allows us to distinguish numerically whether $\varepsilon<\widetilde{\varepsilon}_{m}$, in which case we may use the derivative formula and perform the Newton step, or $\varepsilon>\widetilde{\varepsilon}_{m}$, in which case we just bisect the current interval.
Figure 4. Behavior of $f_{c}(\varepsilon)$ for Example in Section 5.7.
6. Numerical tests
In this section we consider a few numerical experiments illustrating the behavior of our algorithm on a few undirected and directed sparse weighted graphs arising in applications.
All tests are made on a Surface Book 3 computer, under the Matlab version 23.2.0.2365128 (R2023b).
Example 1, matrix $A_{27}$
The considered graph is one of a sequence of graphs arising from molecular dynamics simulations aimed at analyzing the structure of water networks under different temperature and pressure conditions (see [11]). It consists of $710$ nodes (corresponding to oxygen atoms), most of which ($613$) have degree $4$. Its pattern is shown in Figure 5.
Figure 5. Patterns of the adjacency matrix for the graph example $A_{27}$ (left) and for the graph example nopoly from Gaertner collection (right)
In Figure 6 we show the entries of the Perron eigenvector of the original (left picture) and perturbed (right picture), obtained for $m=2$.
Figure 6. Distribution of the entries of Perron eigenvector for the example $A_{27}$. Left picture: original matrix. Right picture: perturbed matrix (for $m=2$).
The five leading entries of the Perron eigenvector of $A$ are the following:
$\mathrm{v}_{26}=0.21705332,\quad\mathrm{v}_{29}=0.24269500,\quad\mathrm{v}_{419}=0.28145420,$
$\mathrm{v}_{111}=0.29422909,\quad\mathrm{v}_{315}=0.37879622.$
After scaling the matrix $A$ to have Frobenius norm $1$, we apply the algorithm.
The norm of the perturbation which determines the coalescence of the two main entries of the Perron eigenvector is $\widetilde{\varepsilon}_{2}=0.0092654$, which is below $1\%$. The two leading entries after the perturbation are
$\mathrm{v}_{111}=0.{\bf 331}01615,\qquad\mathrm{v}_{315}=0.{\bf 331}36357.$
The number of outer iterations is $16$. The CPU time is about 1.23 seconds.
$k$ $\varepsilon_{k}$ $f_{c}(\varepsilon_{k})$#\## eigs
 00 $0.005$ $0.000375176389506$ $25$
$1$ $0.005136626814311$ $0.000351719368954$ $2$
$2$ $0.008469968038737$ $0.000013522265277$ $34$
$3$ $0.008489282220273$ $0.000012884575588$ $2$
$6$ $0.009184673253577$ $0.000000185817345$ $2$
$10$ $0.009212325217551$ $0.000000093285281$ $1$
Table 1. Computation of $\varepsilon_{k}$, $f_{c}(\varepsilon_{k})=F_{\varepsilon}^{c}(E_{k})$ and number of eigenvalue computations in the inner iterations for Example 1, matrix $A_{27}$ with $m=2$.
Note that the relative variations of the two entries corresponding to the nodes with highest centrality are $\delta_{111}=+0.1281$ and $\delta_{315}=-0.1235$.
[…]
$\mathrm{v}_{419}=0.{\bf 3000}0085,\quad\mathrm{v}_{315}=0.{\bf 3000}2148,\quad\mathrm{v}_{111}=0.{\bf 3000}2174.$
[…]
$\mathrm{v}_{109}=0.{\bf 2317}9527,\quad\mathrm{v}_{419}=0.{\bf 2318}0413,\quad\mathrm{v}_{111}=0.{\bf 2318}0684,\quad\mathrm{v}_{315}=0.{\bf 2318}0693.$
The CPU time is about $16$ seconds. It is interesting to note the the node $109$ was not at the highest rank but replaced the node $29$ after the perturbation.
Example $2$, Gaertner matrix
This matrix has been contributed to the TAMU Sparse Matrix Collection by Dr. Klaus Gaertner, Weierstrass Institute, Berlin [^1] See https://sparse.tamu.edu/Gaertner. The matrix, described as the adjacency matrix of a weighted undirected graph, is symmetric. However, some of the weights are negative and there are nonzero entries on the main diagonal. To make it non-negative we consider its absolute value. We also set the diagonal to zero. The corresponding graph consists of $n=10774$ nodes. Its pattern is shown in Figure 5.
The five leading entries of the Perron eigenvector of $A$ are the following:
$\mathrm{v}_{1155}=0.12197226,\quad\mathrm{v}_{3391}=0.12460486,\quad\mathrm{v}_{3002}=0.12460960,$
$\mathrm{v}_{3007}=0.12502517,\quad\mathrm{v}_{3397}=0.12559705.$
After normalizing the matrix $A$ in the Frobenius norm, we apply the algorithm with the aim of obtaining the largest $m$ entries of the Perron eigenvector to match to a $3$-digit accuracy.
First we apply the algorithm with $m=2$. The method starts with $\varepsilon_{0}=10^{-3}$ and sets the tolerance for the stopping criterion to $10^{-6}$. It makes $3$ outer iterates and the final value of the functional is $8\cdot 10^{-12}$. The norm of the perturbation which determines the matching of the $2$ leading entries of the Perron eigenvector is $\widetilde{\varepsilon}_{2}=0.00383633$, which is about $0.00008\%$. The computed $2$ leading entries are the following:
$\mathrm{v}_{3007}=0.{\bf 12558}011,\quad\mathrm{v}_{3397}=0.{\bf 12558}553.$
The CPU time is about $19.2$ seconds.
Then we apply the algorithm with $m=3$. The method starts with $\varepsilon_{0}=10^{-3}$ and sets the tolerance for the stopping criterion to $10^{-10}$. The method makes $8$ outer iterates. The norm of the perturbation which determines the coalescence of the $3$ leading entries of the Perron eigenvector is $\widetilde{\varepsilon}_{3}=0.003909011$, which is about $0.00008\%$. The computed $3$ leading entries are the following.
$\mathrm{v}_{3007}=0.{\bf 12550}017,\quad\mathrm{v}_{3002}=0.{\bf 12550}316,\quad\mathrm{v}_{3397}=0.{\bf 12550}872$
which coincide to a $5$-digit approximation. The CPU time is about $30.8$ seconds.
Finally we set $m=5$. The norm of the perturbation which determines the coalescence of the $5$ leading entries of the Perron eigenvector is $\widetilde{\varepsilon}_{5}=0.0161$, which is about $0.00016\%$. The new leading entries follow:
$\mathrm{v}_{1155}=0.{\bf 124}48751,\quad\mathrm{v}_{3391}=0.{\bf 124}51103,\quad\mathrm{v}_{3007}=0.{\bf 124}53422,$
$\mathrm{v}_{3002}=0.{\bf 124}60202,\quad\mathrm{v}_{3397}=0.{\bf 124}63431.$
The CPU time is about $321.8$ seconds.
Finally, in Figure 7 we show the entries of the Perron eigenvector of the perturbed matrix obtained setting $m=10$.
Figure 7. Distribution of the entries of Perron eigenvector of the matrix obtained for $m=10$ for the example nopoly from Gaertner collection.
7. Global stability of the Perron eigenvector
Here we briefly address problem (II), where we replace the constraint (3) with
$\widetilde{a}_{ij}\geq 0\qquad\mbox{for}\quad(i,j)\in\mathcal{E}.$
According to Theorem 3.1, for perturbations $\widehat{A}$ of $A$ such that the perturbed matrix $\widehat{A}$ remains irreducible, there exists an eigenvalue $\lambda_{1}$ of $\widehat{A}$ which is equal to the spectral radius $\rho(\widehat{A})$. In general, however, it is possible that the second largest eigenvalue $\lambda_{2}$ of $\widehat{A}$ coalesces with $\lambda_{1}$, which would imply that $\widetilde{A}$ becomes reducible. The opposite is not true in general, since $\widetilde{A}$ might become reducible with $\lambda_{1}$ remaining simple.
In the case of coalescence of $\lambda_{1}$ and $\lambda_{2}$, the Perron eigenvector would be no longer stably identified, so that if this occurs due to a small perturbation we consider eigenvector centrality to be poorly stable.
Again, let $\widehat{\mathcal{G}}=(\mathcal{V},\mathcal{E},\widehat{A})$ denote a graph with the same vertices and the same edges of $\mathcal{G}$. We propose - both for undirected and directed graphs - the following stability measure for eigenvector centrality, which we call stability radius:
$\omega(\mathcal{G})=$ $\>\inf_{\widehat{\mathcal{G}}}{\rm dist}\left(\mathcal{G},\widehat{\mathcal{G}}\right)=\inf_{\widehat{A}}\|A-\widehat{A}\|$
under the condition that the Perron eigenvalue of $\widehat{A}$ is not simple and under the constraints $\widehat{a}_{ij}\geq 0$ for $(i,j)\in\mathcal{E}$ and $\widehat{a}_{ij}=0$ for $(i,j)\notin\mathcal{E}$.
The choice of the norm is naturally important; here we use the Frobenius norm for a mainly practical reason: being an inner product norm, we are able to compute the gradient of a suitable functional associated to the problem and proceed in a similar fashion as the one considered in previous section, when we addressed the problem of Section 3.1.
Note that for a strongly connected graph we have the natural lower bound
$\omega(\mathcal{G})<\min\limits_{(i,j)\in\mathcal{E}}a_{ij}.$
In the case of an undirected graph, by the Hoffman–Wielandt theorem [24, p. 368], it is immediate to get the lower bound which relates the stability radius with the spectral gap of $A$ (see [1]).
Theorem 7.1.
Assume $\mathcal{G}$ is an undirected graph to which is associated the weighted adjacency matrix $A$. Then, denoting by $\lambda_{1}\geq\lambda_{2}$, the two largest eigenvalues of $A$, it holds
$\omega(\mathcal{G})\geq(\lambda_{2}-\lambda_{1})/\sqrt{2}.$
Proof.
The proof is based on the observation that if we consider the canonical decomposition of $A=U\Lambda U^{\top}$ and replace $\Lambda$ by $\widehat{\Lambda}$ which differs from $\Lambda$ only in the first two eigenvalues which are set equal to $(\lambda_{1}+\lambda_{2})/2$, one gets a symmetric matrix $\widehat{A}$ with coalescent largest two eigenvalues at a distance $(\lambda_{2}-\lambda_{1})/\sqrt{2}$ from $A$.
Then, considering an arbitrary symmetric matrix $\widetilde{A}$ with coalescent two largest eigenvalues it is straightforward to prove that its distance from $A$ is larger that $(\lambda_{2}-\lambda_{1})/\sqrt{2}$. ∎
If the perturbed matrix $\widehat{A}$ becomes reducible but the Perron eigenvalue remains simple, one may either stop the procedure or continue separately with the two subgraphs.
The stability of the Perron eigenvector is determined by the smallest perturbation of the weighted adjacency matrix $A$ which causes the coaescence of the largest eigenvalue $\lambda_{1}$ with another eigenvalue, This may occur through the coalescence of $\lambda_{1}$ either with a real eigenvalue or with a complex conjugate pair with vanishing imaginary part. We mention that in the special case of irreducible (row) stochastic matrices, the stability of the left Perron vector (stationary probability distribution) under perturbations has been investigated by various authors; see, e.g., [8, 9, 12, 26] and references therein. In these papers the sensitivity of the stationary distribution under perturbations that preserve the irreducibility and stochasticity of the transition probability matrix is expressed in terms of various quantities including the group inverse, the mean passage times, and the so-called ergodicity coefficient.
Also in this case we follow the two-level method considered previously in Section 3.2, to iteratively minimize the functional (for given $\varepsilon$)
(35) $S_{\varepsilon}(E)=\frac{1}{2}\Bigl{(}\lambda_{1}\left(A+\varepsilon E\right)-{\mbox{\rm Re}}\lambda_{2}\left(A+\varepsilon E\right)\Bigr{)}^{2},$
being $\lambda_{2}$ the closest eigenvalue to $\lambda_{1}$ (conventionally with positive imaginary part), under the constraints of unit Frobenius norm of $E$, of non-negativity $A+\varepsilon E\geq 0$ and the preservation of the sparsity pattern of $A$.
In the outer iteration we compute the optimal $\varepsilon$, denoted $\varepsilon_{\star}$,
(36) $\varepsilon_{\star}\longrightarrow\arg\min\limits_{\varepsilon>0}S_{\varepsilon}\left(E(\varepsilon)\right)=0$
by a combined Newton-bisection method. The existence of the minimizer is obtained by compactness arguments. We then have
$\omega(\mathcal{G})=\varepsilon_{\star}.$
Making use of Lemma 4.1, we have the following result:
Lemma 7.1 (Norm unconstrained gradient of the functional).
Assume that the non-negative matrix $A+\varepsilon E$ is irreducible; then, in the space of real matrices with sparsity pattern ${\mathcal{E}}$ we have
$\frac{d}{dt}S_{\varepsilon}\left(E(t)\right)=\varepsilon\Big{\langle}H_{\varepsilon}(E),\dot{E}\Big{\rangle},$
where $H_{\varepsilon}(E)$ has the sparsity pattern determined by the set of edges $\mathcal{E}$ and is given by
(37) $H_{\varepsilon}(E)=P_{\mathcal{E}}\Bigl{(}\kappa_{1}x_{1}y_{1}^{\top}-\kappa_{2}{\mbox{\rm Re}}\left(x_{2}y_{2}^{*}\right)\Bigr{)}$
where $x_{1}$ and $y_{1}$ are the left and right eigenvectors of $A+\varepsilon E$ associated to the Perron eigenvalue $\lambda_{1}$, normalized according to (8), and $x_{2}$ and $y_{2}$ are the left and right eigenvectors associated to $\lambda_{2}$, and $\kappa_{1}=1/(x_{1}^{\top}y_{1}),\kappa_{2}=1/(x_{2}^{*}y_{2})$.
Proof.
The proof makes use of Lemma 4.1 and follows similar steps to the proof of Lemma 5.1. ∎
Note that in the undirected case, i.e. when $A$ and $E$ are symmetric, $\kappa_{1}=\kappa_{2}=1$ and $x_{2}=y_{2}$ is a real vector, so that
$H_{\varepsilon}(E)=P_{\mathcal{E}}\Bigl{(}x_{1}x_{1}^{\top}-x_{2}x_{2}^{\top}\Bigr{)}$
The constrained gradient system
The positivity constraint for $A+\varepsilon E$ can be treated again by penalization, as described in Section 5.5; hence we get the structure and norm constrained gradient system for the penalized functional $S_{\varepsilon}\left(E(t)\right)+\displaystyle\frac{c}{\varepsilon}P_{\varepsilon}\left(E(t)\right)$,
(38) $\dot{E}=-\left(H_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E)\right)+\eta E\quad\hbox{ with }\quad\gamma=\langle E,H_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E)\rangle.$
with $H_{\varepsilon}(E)$ given by (37).
Note that as $\lambda_{2}\to\lambda_{1}$ the terms $\kappa_{1}$ and $\kappa_{2}$ diverge to +$\infty$+ ∞, which implies that we have to use a suitable tolerance to establish the occurrence of coalescence.
7.1. The outer iteration
The outer iteration is analogous to the one considered previously in Section 5.9; in particular, after considering the same penalty function (21), Lemma 5.2 applies verbatim.
8. Conclusions
In this paper we have presented a methodology for finding a minimal perturbation to the adjacency matrix of a weighted (sparse) graph that causes the leading entries of the Perron eigenvector of the perturbed matrix to coalesce. The perturbation is constrained so as to preserve the sparsity of the original matrix. One of the motivations for this work was the question of the reliability and robustness of eigenvector centrality: if a small perturbation causes coalescence of the leading entries of the Perron eigenvector, this centrality measure cannot be considered reliable in the presence of uncertainties or noise, especially when the leading entries of the perturbed eigenvector differ from the original ones. Our methodology easily allows to incorporate additional constraints; for example, only a subset of the graph’s edge weights may be allowed to change. This feature can be used to identify which edges play an important role in influencing the ranking of nodes in the graph. The proposed algorithm appears to perform well on graphs arising from realistic applications, and is fairly efficient in spite of its complexity.